A 6.52 g sample of copper is contaminated with an additional 0.75 g of zinc. Suppose an atomic mass measurement was performed on this sample.

What would be the measured atomic mass?
Answer #1

you need to find the fraction of copper and zinc and multiply it with the mass respectively

hence fraction of copper is 6.52/7.27 and multiply with mass of copper

so (6.52/7.27) * 63.55 = 56.993

in the same way calculatee for zinc (0.75/7.27) * 65.38 = 6.744

so the measured atomic mass is 56.993 + 6.744 = 63.737